Solve:(7x-1)^(1/3)= 3, solve for x
Answer by kapilsinghi(68)

(7x-1)^(1/3)= 3
Cubing Both sides
7x-1 = 27
7x = 27+1
x= 28/7
x = 4
Regards,
Kapil
You can send your problems directely to kapilsinghi123@gmail.com